# Towards Goal-oriented Large Language Model Prompting: A Survey

### v1

@@ Abstract
Large Language Models (LLMs) have shown prominent performance in various downstream tasks in which prompt engineering plays a pivotal role in optimizing LLMs’ performance. This paper, not as an overview of current prompt engineering methods, aims to highlight the limitation of designing prompts while holding an anthropomorphic assumption that expects LLMs to think like humans. From our review of 35 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs. Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework by summarizing ten applicable tasks. With four future directions proposed, we hope to further emphasize and promote goal-oriented prompt engineering.
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT have garnered significant interest for their ability to assimilate vast amounts of information from large corpora of data. They have demonstrated proficiency in, for example, open-domain dialogue Deng et al. (2023); Yu et al. (2023), reasoning Wei et al. (2022); Yao et al. (2023); Besta et al. (2023), planning Song et al. (2023); Liu et al. (2023a), and so on. The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into their potential Liu et al. (2023b). However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al. (2021). Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
@@
To obtain an optimal prompt for tasks, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information. Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996). In interacting with LLMs, which often display human-like conversational abilities, we might be tempted to attribute human-like thought processes to them, expecting them to know from the start that they should decompose complex problems into simpler tasks. This anthropomorphic assumption does not hold for LLMs and can lead to unsatisfactory outcomes Abercrombie et al. (2023). On the other hand, as shown in our literature review of 35 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly. In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers generated from LLMs that result in more effective and reliable human-computer interaction.
@@
In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action implementation, sub-goal result evaluation, and sub-goal selection. We summarize the application of reviewed works on a wide variety of tasks to showcase the broad impact of goal-oriented philosophy in prompt engineering. Finally, we offer four main promising future directions, including the synergy of stages in the framework, applications to other tasks, efficiency problems, and hierarchical decomposition, for further improvement.
@@ 2 Taxonomy of Methods
Figure 1: An overview of the goal-oriented framework for prompting LLMs taking solving a math word problem as an example. (1) Decomposing goal into sub-goal sequences. (2) Action selection for attaining sub-goals. (3) Implementing actions to get sub-goal results. (4) Evaluation of sub-goal achievement based on sub-goal results. (5) Selecting valuable sub-goals for effective goal achievement. Note that stages (2)(3)(4) are taken for all the decomposed sub-goals.
In this section, we categorize existing methods into a goal-oriented framework, as shown in Fig. 1. We further refine the works in each part of the framework according to their distinctive features.
@@ 2.1 Decomposing Goal into Sub-goal Sequences
Decomposing a high-level goal into sub-goals is particularly useful for complex problems where a straightforward answer isn’t sufficient. In this section, we introduce the methods to decompose goals and auxiliary techniques to help improve the performance of decomposition.
@@
Iterative decomposition generates a sub-goal and executes actions to get subgoal results, then repeat this process with the knowledge of the previous sub-goal and sub-goal results. Chain-of-thought prompting (CoT) Wei et al. (2022) can be considered the first work to try to decompose a goal into sub-goal sequences. By showing LLMs a series of intermediate reasoning steps that lead to the final output in the prompt, LLMs can naturally imitate a human-like problem-solving process. Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to the final solution of the problem. Kojima et al. (2022) even found that simply adding “Let’s think step by step.” to the prompt can guide LLMs to perform CoT decomposition as well. The above two works implicitly follow iterative decomposition since LLMs generate tokens in an autoregressive way which means that they can decide the next sub-goal based on previous content.
There are also works that explicitly ask LLMs to follow the iterative decomposition. DecomP Khot et al. (2022) and Successive Prompting Dua et al. (2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks. The background knowledge attained by answering each sub-question are sub-goals to be achieved. Different from CoT which may generate sub-questions sequentially in one output, these two works explicitly repeat prompting LLMs to ask follow-up questions. Self-ask Press et al. (2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions. Thus, LLMs can generate all essential questions and their answers by prompting once. Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions. This finding indicates the significance of decomposing complex goals into simple sub-goals for LLMs.
@@
In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence at once, which means that the latter sub-goal will not be affected by the former ones. For example, Least-to-most prompting Zhou et al. (2022) only prompts LLMs two times, one for generating a plan to decompose the goal into a sub-goal sequence, and the other one for executing the plan. Plan-and-solve prompting Wang et al. (2023a) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output. DEPS Wang et al. (2023c) and GITM Zhu et al. (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft many types of items and tools. In this game, obtaining base materials that are needed to craft a given item can be viewed as sub-goals. While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
@@
The above two categories rely on LLM’s knowledge to decompose the goal into sub-goals. However, due to their hallucination problem Ji et al. (2023), they sometimes generate seemingly plausible sub-goals but not grounded in reality. To ensure the accuracy of decomposition, LLM+P Liu et al. (2023a) and SayPlan Rana et al. (2023) take advantage of classical planners. They use LLMs to translate goals written in natural language to planning domain definition language so that classical planners can deal with them. The output of planners will then be translated back into natural language by LLMs for execution.
Reduce sub-goal space.
Choosing possible sub-goals from a limited space is beneficial for getting a more precise and efficient sub-goal sequence since it avoids LLM getting distracted by irrelevant or even wrong sub-goals. PEARL Sun et al. (2023) is specially designed for question answering over long documents. Several sub-goals like “Finding the definition of A”, “Compare A and B”, and “Summarize A” are pre-set and LLMs only need to choose and arrange valuable sub-goals into a plan. Similar to PEARL, ProCoT Deng et al. (2023) has pre-defined sub-goal sets covering query clarification, topic transition, and negotiation strategy for dialogue systems. DecomP Khot et al. (2022) selects sub-goals from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task. SayPlan Rana et al. (2023) is designed for robot planning. Given a task instruction, semantic search is conducted to identify a relevant subgraph from the whole 3D scene graph that serves as the environment. LLMs will plan based only on this subgraph.
@@ 2.2 Action Selection for Attaining Sub-goals
Action selection is vital for achieving sub-goals as it involves choosing the most effective and valid actions to reach desired outcomes. CoT is a naive way for action selection where it is completely decided by LLMs themselves. However, due to the hallucination problem which is the same in decomposing sub-goals, actions generated by LLMs are often invalid. In this section, apart from naive selection, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
@@
These kinds of methods first set constraints on the action space and then let LLMs decide the actual action. It shares similar ideas with the methods aimed at reducing sub-goal space, which makes it easier for LLMs to select valid actions. MWP Zhang et al. (2023) is proposed for solving math word problems. To ensure the correctness and coherence of the intermediate solution steps, the authors employ an operation prediction module that predicts the needed calculation operation (e.g. multiplication, summation) in this step. LLMs are used to fill the values in this template. In dialogue systems, RLP Hao et al. (2023a) additionally sets a mental state to reflect its personality traits to guide action selection. The authors found that this strategy facilitates contextually rich, coherent, and engaging interactions for LLM-based conversational systems.
@@
Different from constrain-then-select, select-then-mapping first uses LLMs to generate actions solely based on their knowledge and then maps the generated action to the most similar one in the valid action space. Zero-shot planners Huang et al. (2022a), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al. (2022) aim to solve agent planning problems in an interactive environment. In such virtual environments, there are only a few admissible actions that can be applied. However, the actions produced naively by LLMs often cannot be mapped exactly to those executable actions. To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity. When users have access to all the admissible actions, it is suitable to define a set for action mapping.
Reference for selection.
It is empirically evaluated by some works that reference actions are helpful for LLMs to decide on actual actions. Reference action means the successful actions LLMs took when experiencing similar situations in the past. Intuitively, taking reference is similar to human skill improvement through practice. GITM Zhu et al. (2023), created for Minecraft, designs a text-based memory mechanism for LLM to store and retrieve gained useful knowledge during action selection. SALP Gramopadhye and Szafir (2023) collects references by choosing examples that have similar tasks to the given query. The examples are added to the input to autoregressively prompt LLMs. Here we want to discriminate between the reference examples and the in-context examples. In-context examples, which are used in CoT, are used to help LLMs understand and perform tasks Brown et al. (2020). For example, if a prompt includes intermediate steps for a reasoning problem, models will understand that they should also provide intermediate steps. The selection of reference examples in the prompt is stricter than the in-context examples. They should not only be identical in format, which is the criteria for selecting in-context examples, but also be similar in content.
2.3 Implementing Actions to Get Sub-goal Results
@@
In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result. Due to the hallucination problem, the results are sometimes wrong. For example, even if LLMs correctly choose multiplication to solve a math problem, the calculation result is just not correct. Besides, as Khot et al. (2022) claimed, some tasks may not be feasible to solve using only an LLM.
@@
Some works are proposed to leverage external tools to guarantee a precise sub-goal result. When deployed to answer open-domain questions, DecomP Khot et al. (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia. Program-of-Thoughts Chen et al. (2022) first translates reasoning processes into Python codes and then runs them in an interpreter. Toolformer Schick et al. (2023) specifically studied which APIs to call and when to call them when using external tools. In their paper, several tools are adopted, including a calculator for mathematical calculation, a calendar for the awareness of time, a BM25-based Wikipedia search engine for getting comprehensive subject information, a fine-tuned language model (LM) for question answering, and a fine-tuned LM for machine translation. Note that it adopts smaller LMs for specific tasks because authors found that fine-tuned LMs outperform LLMs on some basic functionalities. Following the idea of Toolformer, HuggingGPT Shen et al. (2023) further extends the available LMs to the scale of all LMs on open-sourced machine learning communities like HuggingFace. Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks. Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
2.4 Evaluation of Sub-goal Achievement based on Sub-goal Results
@@
DEPS  Wang et al. (2023c)
VLM DEPS  Wang et al. (2023c)
@@
In CoT, there is no evaluation process for the sub-goal results. The results generated by LLMs are assumed to be correct by default. Similar to action implementation and action selection, sub-goal results suffer from the hallucination problem as well. In addition, mistakes in achieving one sub-goal may accumulate, which can lead to huge deviations from the true answer. Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time. Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well. The types of feedback source and their related works are summarized in Table 1.
Evaluated by LLMs.
Self-refine Madaan et al. (2023) is the first work addressing the evaluation of sub-goal results. After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output. For example, the prompt for providing feedback on code optimization problems may address efficiency and readability. Then, the feedback and initial output are fed together to LLMs for output refinement. Getting feedback and refinement are implemented as an iterative process. During iterations, all of the past feedback and refinement are appended to the prompt. For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified. SelfCheck Miao et al. (2023) decomposes the Self-refine evaluation into finer-grained processes. First, actions and sub-goal results are put to LLMs to summarize their intention. Then, the goal combined with all previous sub-goals, selected actions, and sub-goal results serves as another input to let LLMs extract useful information from them. Finally, only being provided with useful information and summarized intention, LLMs generate the sub-goal result again. This result will be compared with the original result. If the regeneration result supports or contradicts the original sub-goal result, it is likely that the original sub-goal result is correct or incorrect, respectively. In Reflexion Shinn et al. (2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result. The score is considered as feedback.
@@
Figure 2: Illustration of approaches for valuable sub-goal selection. (a) CoT selects all sub-goals in one sub-goal sequence. (b) Self-consistency selects sub-goals based on majority votes. (c) MCR selects sub-goals from multiple sub-goal sequences. (d) Recmind and Selection-Inference select one sub-goal from candidates at each step. (e) ToT explores sub-goals from a tree structure space. (f) GDP-Zero and RAP introduce backpropagation to ToT to balance exploration and exploitation. (g) GoT models all sub-goals in a graph structure space.
Evaluated by external evaluators.
@@
Self-refine and SelfCheck aim to prompt LLMs to stimulate the ability of self-correction. Alternatively, we could seek external evaluators for help to provide feedback, which is similar to the idea of applying external tools for action implementation.
@@
Self-debug Chen et al. (2023) leverages unit tests and executors to gain feedback on generated code snippets. The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection. This finding indicates that external evaluators may be more beneficial than self-reflection thanks to more precise feedback. SayPlan Rana et al. (2023) and Re-prompting Raman et al. (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions. These two methods rely on the error messages that are pre-built into the program of virtual environments. The quality of error messages (e.g. detail) has a significant influence and sometimes we can hardly define error messages in advance. LLM-Planner Song et al. (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed. Inner Monologue Huang et al. (2022b) is a combination of SayPlan and LLM-Planner. It not only gets error messages from environments but also leverages VLMs to attain context information. What’s more, the VLM can also perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information. Similar to Inner Monologue, DEPS Wang et al. (2023c), a method proposed for Minecraft, also gets feedback from both VLMs and environments. DEPS further introduces a step to ask LLMs to explain the reason according to the feedback. GITM Zhu et al. (2023), also used for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation. All of the information can be obtained by calling Minecraft APIs. Reflexion Shinn et al. (2023) proposes various evaluators to handle different tasks. It has reward functions for reasoning tasks based on exact match grading and heuristic functions tailored to specific criteria for decision-making tasks. We refer to them as heuristic rules. Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine. Instead of implementing the action to get the sub-goal result again, these methods restart from action selection.
Reference for evaluation.
Recall that LLM-Planner does not have feedback from the environment. And in Reflexion it can only get a score value indicating the quality of the sub-goal result. For such methods, reference examples play a crucial role in helping LLMs evaluate the sub-goal result with more detail Shinn et al. (2023). The reference selection criteria are the same as in action selection, and examples with similar actions are more beneficial Song et al. (2023). For example, “cook a potato” is likely more informative than “clean a plate” for a sub-goal result from “cooking an egg”. While Reflexion builds a small memory buffer to store recent trials, LLM-Planner selects reference examples from the whole training sets based on the Euclidean distance between the BERT embeddings of references and the running trial.
2.5 Selecting Valuable Sub-goals for Effective Goal Achievement

### v2

# Towards Goal-oriented Prompt Engineering for Large Language Models: A Survey

@@ Abstract
Large Language Models (LLMs) have shown prominent performance in various downstream tasks and prompt engineering plays a pivotal role in optimizing LLMs’ performance. This paper, not only as an overview of current prompt engineering methods, but also aims to highlight the limitation of designing prompts based on an anthropomorphic assumption that expects LLMs to think like humans. From our review of 36 representative studies, we demonstrate that a goal-oriented prompt formulation, which guides LLMs to follow established human logical thinking, significantly improves the performance of LLMs. Furthermore, We introduce a novel taxonomy that categorizes goal-oriented prompting methods into five interconnected stages and we demonstrate the broad applicability of our framework. With four future directions proposed, we hope to further emphasize the power and potential of goal-oriented prompt engineering in all fields.
Towards Goal-oriented Prompt Engineering
for Large Language Models: A Survey
Haochen Li    Jonathan Leung    Zhiqi Shen Nanyang Technological University, Singapore {haochen003, jonathan008, zqshen}@ntu.edu.sg
@@ 1 Introduction
Large Language Models (LLMs) have garnered significant interest with their ability to assimilate vast amounts of information from large corpora of data. They have demonstrated proficiency in open-domain dialogue Deng et al. (2023); Yu et al. (2023), reasoning Wei et al. (2022); Yao et al. (2023); Besta et al. (2023), planning Song et al. (2023); Liu et al. (2023a), and many more tasks. The predominant approach to leveraging LLMs for downstream tasks involves crafting tailored text prompts to tap into LLMs’ potential Liu et al. (2023c). However, the effectiveness of LLMs is contingent upon the prompting strategy employed, leading to performance disparities Zhao et al. (2021). Such a phenomenon has given rise to the field of prompt engineering, which investigates the optimal formulation of prompts for specific tasks.
@@
To prompt LLMs appropriately, it is important to first understand the deviation in the human way of thinking and LLMs’ way of processing information. Humans are taught by experience to approach complex objectives by breaking down the main goal into more manageable sub-goals, a strategy supported by goal-setting theory Austin and Vancouver (1996). In interacting with LLMs, which often display human-like conversational abilities, we might be tempted to attribute human-like thought processes to them, expecting them to know from the start that they should decompose complex problems into simpler tasks. This anthropomorphic assumption does not hold for LLMs and can lead to unsatisfactory outcomes Abercrombie et al. (2023). On the other hand, as shown in our literature review of 36 representative studies, if we design the prompts from a goal-oriented perspective that guides LLMs to mimic human thinking, LLMs’ performance improves significantly. In the meantime, aligning LLM’s behavior with human logic allows more human-interpretable answers from LLMs and results in more effective and reliable human-computer interaction.
Figure 1: An overview of the goal-oriented framework for prompting LLMs taking solving a math word problem as an example. (1) Decomposing goal into sub-goal sequences. (2) Action selection for attaining sub-goals. (3) Executing actions to get sub-goal results. (4) Evaluating sub-goal results. (5) Further selection of valuable sub-goals. Note that stages (2)(3)(4) are taken for all the decomposed sub-goals.
While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures. Some surveys summarize studies that focus on specific tasks or problems. Qiao et al. (2022); Huang and Chang (2022) focus solely on reasoning tasks and Wang et al. (2024) on planning tasks. Huang et al. (2023) deals with the hallucination problem of LLMs. Other studies focus on generally applicable prompting but different sub-fields. Amatriain (2024); Sahoo et al. (2024) simply listed out recent techniques without systematic categorization. Liu et al. (2023c) focus on the prompting framework for facilitating the interaction between LLMs with external worlds. Liu et al. (2023b) focuses on methods for learnable continuous prompt, while our work studies discrete prompt. As many of the will-functioned LLMs are not open-sourced (e.g. GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
@@
In this paper, we propose the first goal-oriented taxonomy of prompting methods, which classifies goal-oriented prompting methods into goal decomposition, action selection, action execution, sub-goal result evaluation, and sub-goal selection, as shown in Fig. 1. We further classify the surveyed methods by their targeted task and analyze their respective performance to provide readers with a better estimate of the current progress of goal-oriented prompt engineering. Finally, we offer four promising future directions, including hierarchical decomposition, efficiency problems, the synergy of stages, and broader applications.
@@ 2 Taxonomy of Methods
In this section, we go into detail about the five steps described in our goal-oriented framework in Fig. 1.
@@ 2.1 Decomposing Goal into Sub-goal Sequences
Decomposing a high-level goal into sub-goals is particularly useful for complex problems where a straightforward answer isn’t sufficient. In this section, we introduce the methods to decompose goals and improve the performance of decomposition.
@@
Iterative decomposition generates a sub-goal, gets sub-goal results, and repeats this process with the knowledge of the previous sub-goal and sub-goal results. Chain-of-thought prompting (CoT) Wei et al. (2022) can be considered the first work that follows iterative decomposition. With either “Let’s think step by step.” Kojima et al. (2022), or intermediate reasoning steps in in-context examples in the prompt, LLMs can imitate the decomposition process and solve the problem step by step. Here, the intermediate reasoning steps can be considered as sub-goals, and they are sequentially connected to form a sub-goal sequence that leads to a correct output. CoT implicitly follows iterative decomposition as LLMs generate tokens in an autoregressive way, with the next sub-goal decided based on previous content.
There are also works that explicitly ask LLMs to follow the iterative decomposition. DecomP Khot et al. (2022) and Successive Prompting Dua et al. (2022) are two contemporary works that repeatedly ask sub-questions to get background knowledge for question answering tasks. The background knowledge required to answer the question is sub-goal to be achieved. Different from CoT which may generate sub-questions sequentially in one output, these two works repeatedly prompt LLMs to ask follow-up questions. Self-ask Press et al. (2022) improves the efficiency of DecomP and Successive prompting by designing a template for follow-up questions. With that template, LLMs can generate all essential questions and their answers by prompting once. Besides, in this paper, the authors empirically show that LLMs are often wrong when asking them to answer a complex question even if they know the true answer of all needed sub-questions. This finding indicates the significance of decomposing complex goals into simple sub-goals for LLMs.
@@
In contrast to iterative decomposition, plan-then-execute decomposition methods generate the sub-goal sequence all at once, which means that the latter sub-goal will not be affected by the result of the former ones. For example, Least-to-most prompting Zhou et al. (2022) only prompts LLMs two times, one for generating a plan to decompose the goal into a sub-goal sequence, and the other one for executing the plan. Plan-and-solve prompting Wang et al. (2023b) further improves the efficiency of Least-to-most prompting by merging the generation of the plan and execution of the plan into one output. DEPS Wang et al. (2023d) and GITM Zhu et al. (2023) are decomposition methods designed for Minecraft, an open-world game in which an agent can craft different items and tools. In this game, obtaining base materials that are needed to craft a target item can be viewed as sub-goals. While DEPS generates a plan to obtain required objects in sequence solely based on LLMs, GITM leverages a pre-defined sub-goal tree to help LLMs locate prerequisites more precisely.
@@
The above two categories rely on LLM’s knowledge to decompose the goal into sub-goals. However, due to the hallucination problem Ji et al. (2023), LLMs sometimes generate seemingly plausible sub-goals but not grounded in reality. To ensure the accuracy of decomposition, LLM+P Liu et al. (2023a) and SayPlan Rana et al. (2023) take advantage of classical planners. They use LLMs to translate goals written in natural language to planning domain definition language so that classical planners can deal with them. The output of planners will then be translated back into natural language by LLMs for execution.
@@ 2.2 Action Selection for Attaining Sub-goals
With sub-goals defined, another important step is action selection, choosing the most effective and valid actions to reach desired outcomes. CoT is a naive way for action selection where it is completely decided by LLMs themselves. However, due to the hallucination problem, actions generated by LLMs are often invalid. In this section, we classify the advanced action selection methods into two classes, constrain-then-select and select-then-mapping.
@@
Constrain-then-select method predefines an action space and then has LLMs select the actual action among the space. As in MWP Zhang et al. (2023), to deal with math word problems, the authors first employ an operation prediction module to predict the needed calculation operation (e.g. multiplication, summation). LLMs are then used to select the appropriate operands and complete the calculation. PEARL Sun et al. (2023) utilizes constrained action space for question answering over long documents. The authors predefine actions like “Finding the definition of A”, “Compare A and B”, and “Summarize A”. DecomP Khot et al. (2022) selects actions from a set like “split” and “merge” for k-th letter concatenation, a symbolic reasoning task. SayPlan Rana et al. (2023) is for robot planning. Given a task instruction, the authors apply semantic search to first identify a task-relevant subgraph from the whole 3D scene graph, then have LLMs plan based on the subgraph. In dialogue systems, RLP Fischer (2023) allows the user to set a mental state of LLM, such as beliefs or desires, to instruct action selection. Such a setting improves the context richness, coherence, and interactivity of LLM-based conversational systems. Similar to PEARL, ProCoT Deng et al. (2023) forms pre-defined action sets covering query clarification, topic transition, and negotiation strategy for dialogue systems.
@@
Different from constrain-then-select, select-then-mapping first uses LLMs to generate actions solely based on their knowledge and then maps the generated action to the most similar one in the valid action space. Zero-shot planners Huang et al. (2022a), SALP Gramopadhye and Szafir (2023), and Re-Prompting Raman et al. (2022) aim to solve agent planning problems in an interactive environment. In such virtual environments, there are only a few admissible actions that can be applied. However, the actions produced naively by LLMs often cannot be mapped exactly to those executable actions. To remedy this, researchers employ a text-embedding language model to translate LLM-generated actions into the most similar admissible actions by calculating cosine similarity.
2.3 Executing Actions to Get Sub-goal Results
@@
In CoT, LLMs rely on their knowledge for action execution to get the sub-goal result. Due to the hallucination problem, the results are sometimes wrong. For example, even if LLMs correctly select multiplication as the action to solve a math problem, the calculation result can be wrong. Besides, as Khot et al. (2022) claimed, some tasks may not be feasible to solve using only one LLM. With such consideration, some works leverage external tools to guarantee a precise sub-goal result.
Some studies specifically study one task Khot et al. (2022); Chen et al. (2022); Wang et al. (2023c). To answer open-domain questions, DecomP Khot et al. (2022) applies an ElasticSearch-based retrieval system to retrieve knowledge from certain knowledge bases like Wikipedia. Program-of-Thoughts (PoT) Chen et al. (2022) first translates reasoning processes into Python codes and then runs them in an interpreter. Recmind Wang et al. (2023c) uses SQL to search for the history of a user in the database in recommender systems.
@@
There are additional studies that aim for multi-tasking. Toolformer Schick et al. (2023) specifically studied which APIs to call and when to call them when using external tools. In their paper, several tools are adopted, including a calculator, a calendar, a BM25-based Wikipedia search engine, and a fine-tuned language model (LM). The authors pointed out that small LMs finetuned for specific tasks can outperform some LLMs on certain task. Following the idea of Toolformer, HuggingGPT Shen et al. (2023) further extends the available LMs to the scale of all LMs on open-sourced machine learning communities like HuggingFace. Specifically, in their paper, they evaluate HuggingGPT on nine types of NLP tasks, nine types of CV tasks, four types of audio tasks, and two types of video tasks. Additionally, with the development of open-sourced communities, this work demonstrates the potential of achieving all the sub-goals that can be described in language by leveraging external tools.
2.4 Evaluating Sub-goal Results
@@
DEPS  Wang et al. (2023d)
VLM DEPS  Wang et al. (2023d)
@@
In CoT, there is no evaluation process for the sub-goal results. The results generated by LLMs are assumed to be correct by default. Answering by LLMs, sub-goal results also suffer from the hallucination problem. In addition, errors may propagate along the pipeline and result in huge deviations from the correct answer. Thus, it is necessary to evaluate the sub-goal results at each step and correct them in time. Note that the methods we review in this section can be applied to evaluate the overall goal achievement as well. The types of feedback source and their related works are summarized in Table 1.
@@
Figure 2: Illustration of approaches for valuable sub-goal selection. (a) CoT selects all sub-goals in one sub-goal sequence. (b) Self-consistency, a variant of CoT, selects sub-goals based on majority votes. (c) MCR, a variant of CoT, selects sub-goals from multiple sub-goal sequences. (d) Recmind and Selection-Inference, variants of CoT, select one sub-goal from candidates at each step. (e) ToT explores sub-goals from a tree structure space. (f) GDP-Zero and RAP, variants of ToT, introduce backpropagation to ToT to balance exploration and exploitation. (g) GoT models all sub-goals in a graph structure space.
Evaluate with LLMs.
Self-refine Madaan et al. (2023) is the first work addressing the evaluation of sub-goal results. After getting the output of LLMs, Self-refine feeds the output combined with task-specific prompts to the same LLM to get feedback for the output. Then, the feedback and initial output are fed together to LLMs for output refinement. The refinement process iterates with all of the past feedback and refinement appended to the prompt. For the correction process of the following works in this section, they all follow Self-refine to correct sub-goal results, if not specified. SelfCheck Miao et al. (2023) decomposes the Self-refine evaluation into finer-grained processes. First, actions and sub-goal results are put to LLMs to summarize their intention. Then, the goal combined with all previous sub-goals, selected actions, and sub-goal results serves as another input to let LLMs extract useful information from them. Finally, only being provided with useful information and summarized intention, LLMs generate the sub-goal result again. This result will be compared with the original result. If the regeneration result supports the original sub-goal result, the original result is considered correct, and vice versa. In Reflexion Shinn et al. (2023), the authors try using another instantiation of an LLM to obtain a score for the sub-goal result.
Evaluate with external evaluators.
@@
In contrast to self-evaluation, we could seek the help of external evaluators for providing feedback, which is similar to but different from the idea of applying external tools for action execution.
@@
Some methods Chen et al. (2023); Wang et al. (2023a); Rana et al. (2023); Raman et al. (2022) leverage the error messages that are pre-built into either the programming languages or the program of virtual environments to evaluate sub-goal results. Self-debug Chen et al. (2023) and INTERVENOR Wang et al. (2023a) leverages unit tests and executors to gain feedback on generated code snippets. The authors also empirically show that leveraging unit test error messages leads to superior performance than LLM’s self-reflection. This finding indicates that external evaluators may be more beneficial than self-reflection thanks to the more precise feedback. SayPlan Rana et al. (2023) and Re-prompting Raman et al. (2022) directly leverage the signs of success or failure with error messages from the environment as feedback to revise their actions. GITM Zhu et al. (2023), a method proposed for Minecraft, designs a structured feedback template to get detailed knowledge of the current situation. All of the information can be obtained from Minecraft API returns.
The quality of error messages (e.g. detail) is influential to the evaluating process, but sometimes it is hard to define error messages in advance. To further address it, LLM-Planner Song et al. (2023) explores the possibility of using Vision-Language Models (VLMs) to provide feedback by describing the environment after confirming that the action has failed. Inner Monologue Huang et al. (2022b) is a combination of SayPlan and LLM-Planner, where VLMs additionally perform the task of visual question answering, which enables a finer-grained human-computer interaction to get valuable information. Similar to Inner Monologue, DEPS Wang et al. (2023d), designed for Minecraft, also gets feedback from both VLMs and environments. DEPS further introduces a step to ask LLMs to explain the reason according to the feedback. Reflexion Shinn et al. (2023) proposes various evaluators to handle different tasks. It has reward functions for reasoning tasks based on exact match grading and heuristic functions tailored to specific criteria for decision-making tasks. We refer to them as heuristic rules. Note that the correction process of SayPlan, Re-prompting, LLM-Planner, Inner Monologue, DEPS, and GITM is different from Self-refine. Instead of executing the action to get the sub-goal result again, these methods restart from action selection.
2.5 Further Selection of Valuable Sub-goals